Suppose the pdf of the magnitude of X of a dynamic load on a bridge is given by ��={ 1 8(1+3�,0≤�ɤ2 0, �ѡℎ�џ�і�ђ i) Determine �0.5≤�ɤ1.5) ii) Obtain the cumulative density function of X

1
Expert's answer
2021-11-29T13:11:41-0500
"f(x)=\begin{matrix}
 1/8+(3/8)x & 0\leq x\leq 2 \\
\\
 0 & otherwise
\end{matrix}"

i)


"P(0.5\leq x\leq 1.5)=\displaystyle\int_{0.5}^{1.5}(1/8+(3/8)x)dx"

"=\bigg[(1/8)x+(3/16)x^2\bigg]\begin{matrix}
 1.5 \\
 0.5
\end{matrix}"

"=\dfrac{1.5}{8}+\dfrac{3(1.5)^2}{16}-\dfrac{0.5}{8}-\dfrac{3(0.5)^2}{16}=0.5"

ii)

"F(x)=\displaystyle\int_{-\infin}^{x}f(x)dx"

"F(x) = \begin{cases}
 0 &x<0 \\
 (1/8)x+(3/16)x^2 &0\leq x<2 \\
 1 & x\geq 2
\end{cases}"